One lens has a focal length of +15.6 cm. A second lens of focal length +21.1 cm is placed in contact with the first lens. What is the equivalent focal length of the combination of lenses?
Answer #1

1/f = 1/f1 + 1/f2

f = (f1xf2)/(f1+f2) = (15.6x21.1)/(15.6+21.1) = 8.96cm